# The art of modeling nuclear reactions with weakly bound nuclei: status and perspectives

### v1

@@ 4.3 Simultaneous inclusion of projectile breakup and target excitation in transfer reactions
The models discusses in the previous subsections are based on a strict three- or four-body description of the transfer reaction in which possible excitations of the target ($A$) are only accounted for approximately by the choice of the effective interactions. From the analysis of reactions with ordinary nuclei, it is known that an explicit account of collective excitations of the colliding nuclei is required for an accurate description of the reaction. For $(d,p)$ reactions on deformed targets, one may anticipate that both projectile breakup and target excitation can play a role and may require their explicit inclusion. This is not possible in the standard formulations of the aforementioned CCBA and ADWA methods, which tend to emphasize one of the two mechanisms. As a way of including the two mechanisms simultaneously, one may resort to the extended CDCC wavefunction (24) which was devised to include in the model space the projectile breakup channel as well as some excited states of the target nucleus.
@@
$\Psi^{(+)}_{d}({\vec{r}},{\vec{R}},\xi)$ $=\phi_{d}({\vec{r}})\Phi^{0}_{A}(\xi)\chi^{(+)}_{d,0}({\vec{R}})+\phi_{d}({\vec{r}})\Phi^{0}_{A}(\xi)\chi^{(+)}_{d,1}({\vec{R}})$
$+\sum_{n}\phi^{n}_{pn}({\vec{r}})\Phi^{\text{1}}_{A}(\xi)\chi^{(+)}_{n,0}({\vec{R}})+\sum_{n}\phi^{n}_{pn}({\vec{r}})\Phi^{\text{1}}_{A}(\xi)\chi^{(+)}_{n,1}({\vec{R}})$ (64)
where $\{\phi_{d},\phi^{n}_{pn}\}$ denote the deuteron ground state and (discretized) continuum states, and $\{\chi^{(+)}_{n,0}({\vec{R}}),\chi^{(+)}_{n,1}({\vec{R}})\}$ the functions describing the projectile-target relative motion with the target in either its ground state $\Phi^{0}_{A}(\xi)$ or in the excited state $\Phi^{\text{1}}_{A}(\xi)$, respectively. Therefore, the first two terms of (4.3) describe, respectively, elastic and inelastic scattering with the deuteron remaining in its ground state. The third and fourth terms describe deuteron breakup with respect to the target in its ground state or first excited state, respectively. When inserted into Eq. (55) this gives rise also to four terms,
@@
As an example, we consider the reaction 10Be(d,p)11Be(g.s.), including both the 10Be(g.s.) and its first excited state ($2^{+}$; $E_{x}=3.367$ MeV). The different reaction paths described above are schematically shown in Fig. 30.
Figure 30: (Color online) Pictorial representation of the 10Be(d,p)11Be reaction, highlighting the different paths considered in this work. The initial $d+$10Be system includes breakup of the deuteron and excitation of 10Be coupled to all orders. The transfer channel is then coupled in first order to these channels, and four paths are distinguished: I (elastic transfer), II (inelastic transfer), III (breakup transfer) and IV (inelastic breakup transfer). Figure taken from Gom17b , with authorization from APS.
Calculations for this reaction have been reported in Ref. Gom17b . The structure of the 11Be nucleus is treated within the particle-rotor model of Ref. Nun96 , which assumes a deformation length of $\delta_{2}=1.664$ fm for 10Be and includes the ground state ($0^{+}$) and first excited state ($2^{+}$) of this nucleus.
In Fig. 31 we show the calculated angular distributions for the deuteron incident energies of $E_{d}=20$ MeV (top) and $E_{d}=60$ MeV (bottom). The different lines shown in each panel correspond to the paths of Fig. 30. In each panel, we show the full calculation (thick solid black line) along with the calculations keeping only one of the terms in Eq. (65). The red solid line corresponds to ${\cal T}^{el}_{dp}$ (path I in Fig. 31), the blue solid only to ${\cal T}^{inel}_{dp}$ (II), the red dashed only to ${\cal T}^{elbu}_{dp}$ (III) and the blue dashed only to ${\cal T}^{inbu}_{dp}$ (IV). As before, for the $d$+10Be partition all channels are coupled to all orders, while the transfer channel is coupled in first order. It can be seen that for both energies the main contributor is the elastic transfer ${\cal T}^{el}_{dp}$, while, at small angles, for $E_{d}=20$ MeV the second main contributor is the breakdown of the deuteron, ${\cal T}^{elbu}_{dp}$, whereas at 60 MeV it is the excitation of 10Be, ${\cal T}^{inel}_{dp}$
@@
In Fig. 32 we plot, as a function of incident energy, the ratio of the cross section at the peak for the calculation where one of the transfer paths was selected divided by the cross section at the peak for the full calculation, containing all transfer paths. As can be seen in the figure, path II, corresponding to ${\cal T}^{inel}_{dp}$ gains relevance as the energy increases,
Figure 32: Ratio between cross sections at the peak as a function of the deuteron incident energy. For each energy the ratio is between the calculation where one of the transfer paths is selected (as in Fig. 31) and the full calculation. As can be seen, path II, corresponding to the inelastic excitation of 10Be, gains importance with increasing energy so that at the higher energies, where the $R_{x}$ factors are smaller, it is of the same magnitude as path III, corresponding to deuteron breakup. Figure taken from Gom17b , with authorization from APS.
@@ 5.1 The Ichimura-Austern-Vincent (IAV) model
This expression has the appealing feature of connecting the NEB cross section with the CDCC three-body wavefunction. As noted in Sect. 3.7, the CDCC does not provide detailed NEB breakup cross sections by itself. However, with the aid of Eqs. (67) and (66) one can effectively extract the NEB content of the CDCC wavefunction.
@@
With this approximation, the NEB component of the $b$ singles cross section becomes
@@
Figure 36: IAV calculations for the reaction 197Au(9Be,8Be*)198Au populating bound states of the residual nucleus. (a) Norm of the overlaps between the 9Be g.s. and the 8Be continuum pseudostates for different relative angular momenta $L$, as a function of the $\alpha$-$\alpha$ relative energy up to 20 MeV. Dashed lines are included as a guide. (b) Differential cross section, at $E_{\mathrm{cm}}=36$ MeV, as a function of the $\alpha$-$\alpha$ relative energy in the final 8Be system, for 0+ states (solid line) and $2^{+}$ states (dashed line). The symbols indicate the nominal energies of the g.s. ($0^{+}$) and the first $2^{+}_{1}$ resonance of 8Be. (c) Angle-integrated total transfer cross section as a function of $E_{\mathrm{cm}}$. The contribution from the $0^{+}$ (dot-dashed line) and $2^{+}$ (dotted) states of 8Be, together with their sum (blue solid), are compared to the experimental data of Ref. gollan21 . The IAV result rescaled by VMC spectroscopic factors (dashed line) is also shown (see text). Adapted from Ref. Vil24 .
@@ 5.2 The Eikonal Hussein-McVoy formula (EHM)
    This particular choice has the virtue of taking into account breakup effects in the entrance channel. However, since each potential in (79) is evaluated in the corresponding fragment-target coordinate, the associated initial state wavefunction of the system would be a solution of a complicated three-body equation.
@@
With the particular choice (79) and the assumption (80) one obtains the following result for the $x$-channel wavefunction:
@@ 5.3 Interpretation of inclusive breakup data
In the previous sections, we have considered the elastic breakup and nonelastic breakup (with the latter possibly including transfer and incomplete fusion) taking place in a breakup reaction. In actual inclusive breakup experiments these contributions will appear entangled in the data, although they produce some distinctive features. For example, the energy distribution of the observed fragment at a given scattering angle will exhibit a characteristic shape, as shown in Fig. 37 for a hypothetical $A(d,pX)$ reaction. To understand this spectrum, it is important to recall that a given proton energy and angle will univocally determine the excitation energy of the residual $n+A$ system. According to the proton energy (or, equivalently, the residual system excitation energy), we may distinguish the following regions:
Figure 37: Proton energy spectrum from a $A(d,pX)$ inclusive breakup reaction.
@@
    As the excitation energy of the residual nucleus increases, so does the density of states and the low-lying bound states will merge to form a quasi-continuum. Here, the treatment with the aforementioned methods becomes more troublesome because of the impossibility of disentangling unambiguously the contribution of each state from the data. When one is interested in energy averaged cross sections, the combination of the IAV model of Sec 5.1 with the dispersive models provides an appealing alternative to the traditional methods.
@@
    As the excitation energy of the residual nucleus increases above the threshold, the narrow resonances disappear, the spectrum becomes continuous and structureless, giving rise to a bell-shaped bump. The most probable energy of the emerging proton can be obtained by assuming that the proton gets half of the kinetic of the deuteron at the point of breakup, plus the Coulomb energy of the deuteron at that place, and half the internal energy of the deuteron ($E_{B}=-2.22$ MeV) (see Ref. Bla91 , p.509), i.e.,
@@
    In the former contributions, we have implicitly assumed that the observed particle (proton in this case) scatters elastically by the target nucleus. There will be situations in which this this not the case; for example, when the projectile fuses completely with the target nucleus forming a compound nucleus that will eventually thermalize by emitting particles and gamma rays. Among these particles, there will be protons that will add up incoherently to the proton spectrum. These protons are typically emitted with low energy in the c.m. frame and will therefore contribute to the low energy part of the proton spectrum. Some phenomena related to the fusion of weakly-bound nuclei will be discussed in the next section.
@@ 6 Fusion involving weakly bound nuclei
The previous sections have been devoted to the modeling of direct nuclear reactions. Fusion between weakly-bound nuclei display also distinctive features which require adequate reaction formalisms and considerations with respect to the case of well-bound nuclei. The topic of fusion with both stable and unstable nuclei has motivated many works and excellent review papers have been published in recent years, so we refer the reader to these works for a detailed account of the present status of the description of fusion (e.g. Can06 ; Can15 ; Can21 ). Here, we focus on two phenomena which are the focus of many theoretical and experimental efforts by several groups. One is the phenomenon of complete fusion suppression. Complete fusion is conveniently defined as the process in which the whole charge of the projectile and target nuclei merge, giving rise to an excited compound nucleus that will subsequently decay by particle and/or gamma emission. The other phenomenon is the large observed yields compatible with the partial fusion of the projectile (incomplete fusion, ICF). Experimental results and theoretical calculations indicate that these two phenomena are actually related since they appear simultaneously and so a plausible explanation of the CF suppression might be in fact the leak of flux going to the ICF channels.
@@
Early analyses of these experiments tried to explain the phenomenon using coupled-channels calculations, including the coupling to low-lying excited states of the projectile and target Das99 ; Das02 ; Kum12 ; Zha14 ; Fan15 . Yet, these calculations systematically failed to reproduce such suppression. This failure has been attributed to the omission in these calculations of the breakup of the projectile; a scenario was suggested in which the weakly bound projectile breaks up prior to reaching the fusion barrier, with the subsequent reduction of probability of complete capture. This interpretation is supported by the presence of large $\alpha$ yields as well as target-like residues which are consistent with the capture of one of the fragment constituents of the projectile, that is, ICF. To account for these observations, some authors have proposed a two-step scenario Das02 ; Dia07 consisting on the elastic dissociation of the projectile followed by the capture of one of the fragments by the target. However, calculations based on a three-dimensional classical dynamical model Dia07 , which incorporates this two-step breakup-fusion mechanism, can only explain a small fraction of the observed CF suppression for 9Be Coo16 and 8Li Coo18 reactions. More encouraging results have been obtained with different methods based on the CDCC formalism, as described in the following subsections.

### v2

@@ 4.3 Simultaneous inclusion of projectile breakup and target excitation in transfer reactions
The models discussed in the previous subsections are based on a strict three- or four-body description of the transfer reaction in which possible excitations of the target ($A$) are only accounted for approximately by the choice of the effective interactions. From the analysis of reactions with ordinary nuclei, it is known that an explicit account of collective excitations of the colliding nuclei is required for an accurate description of the reaction. For $(d,p)$ reactions on deformed targets, one may anticipate that both projectile breakup and target excitation can play a role and may require their explicit inclusion. This is not possible in the standard formulations of the aforementioned CCBA and ADWA methods, which tend to emphasize one of the two mechanisms. As a way of including the two mechanisms simultaneously, one may resort to the extended CDCC wavefunction (29) which was devised to include in the model space the projectile breakup channel as well as some excited states of the target nucleus.
@@
$\Psi^{(+)}_{d}({\vec{r}},{\vec{R}},\xi)$ $=\phi_{d}({\vec{r}})\Phi^{0}_{A}(\xi)\chi^{(+)}_{d,0}({\vec{R}})+\phi_{d}({\vec{r}})\Phi^{1}_{A}(\xi)\chi^{(+)}_{d,1}({\vec{R}})$
$+\sum_{n}\phi^{n}_{pn}({\vec{r}})\Phi^{\text{0}}_{A}(\xi)\chi^{(+)}_{n,0}({\vec{R}})+\sum_{n}\phi^{n}_{pn}({\vec{r}})\Phi^{\text{1}}_{A}(\xi)\chi^{(+)}_{n,1}({\vec{R}})$ (72)
where $\{\phi_{d},\phi^{n}_{pn}\}$ denote the deuteron ground state and (discretized) continuum states, and $\{\chi^{(+)}_{n,0}({\vec{R}}),\chi^{(+)}_{n,1}({\vec{R}})\}$ the functions describing the projectile-target relative motion with the target in either its ground state $\Phi^{0}_{A}(\xi)$ or in the excited state $\Phi^{\text{1}}_{A}(\xi)$, respectively. Therefore, the first two terms of Eq. (4.3) describe, respectively, elastic and inelastic scattering with the deuteron remaining in its ground state. The third and fourth terms describe deuteron breakup with respect to the target in its ground state or first excited state, respectively. When inserted into Eq. (63) this gives rise also to four terms,
@@
As an example, we consider the reaction 10Be(d,p)11Be(g.s.), including both the 10Be(g.s.) and its first excited state ($2^{+}$; $E_{x}=3.367$ MeV).
Calculations for this reaction have been reported in Ref. Gom17b . The structure of the 11Be nucleus is treated within the particle-rotor model of Ref. Nun96 , which assumes a deformation length of $\delta_{2}=1.664$ fm for 10Be which allow for the coupling of the ground ($0^{+}$) and first excited state ($2^{+}$) of this nucleus.
In Fig. 34 we show the calculated angular distributions for the deuteron incident energies of $E_{d}=20$ MeV (top) and $E_{d}=60$ MeV (bottom). In each panel, we show the full calculation (thick solid black line) along with the calculations keeping only one of the terms in Eq. (73). The red solid line corresponds to ${\cal T}^{el}_{dp}$ (path I in Fig. 34), the blue solid to ${\cal T}^{inel}_{dp}$ (II), the red dashed to ${\cal T}^{elbu}_{dp}$ (III) and the blue dashed to ${\cal T}^{inbu}_{dp}$ (IV). As before, for the $d$+10Be partition all channels are coupled to all orders, while the transfer channel is coupled in first order. This is shown in a pictorial representation in Fig. 35, where the black lines with two arrows correspond to couplings to all orders, while the red and blue lines with one arrow denote couplings to first order. The labels related to each of the terms in Eq. 73 are also included. It can be seen that for both energies the main contributor is the elastic transfer ${\cal T}^{el}_{dp}$, while, at small angles, for $E_{d}=20$ MeV the second main contributor is the breakup of the deuteron, ${\cal T}^{elbu}_{dp}$, whereas at 60 MeV it is the excitation of 10Be, ${\cal T}^{inel}_{dp}$
@@
Figure 35: (Color online) Pictorial representation of the 10Be(d,p)11Be reaction, highlighting the different contributions considered in this work. The initial $d+$10Be system includes breakup of the deuteron and excitation of 10Be coupled to all orders (black lines). The transfer channel is then coupled in first order to these channels (red and blue lines), and four terms are distinguished: I (elastic transfer), II (inelastic transfer), III (breakup transfer) and IV (inelastic breakup transfer). Figure adapted from Gom17b , with authorization from APS.
In Fig. 36 we plot, as a function of incident energy, the ratio of the cross section at the peak for the calculation where one of the transfer paths was selected divided by the cross section at the peak for the full calculation, containing all transfer paths. As can be seen in the figure, path II, corresponding to ${\cal T}^{inel}_{dp}$ gains relevance as the energy increases, although it must be noted that this behaviour is found for this particular reaction and should not be generalized to other target nuclei.
Figure 36: Ratio between cross sections at the peak as a function of the deuteron incident energy. For each energy the ratio is between the calculation where one of the transfer terms is selected (as in Fig. 34) and the full calculation. As can be seen, term II, corresponding to the inelastic excitation of 10Be, gains importance with increasing energy so that at the higher energies, it is of the same magnitude as term III, corresponding to deuteron breakup. Figure taken from Gom17b , with authorization from APS.
@@ 5.1 The Ichimura-Austern-Vincent (IAV) model
This expression has the appealing feature of connecting the NEB cross section with the CDCC three-body wavefunction. As noted in Sec. 3.8, the CDCC method does not provide detailed NEB breakup cross sections by itself. However, with the aid of Eqs. (75) and (74) one can effectively extract the NEB content of the CDCC wavefunction.
@@
where $\phi^{0}_{A}$ is the target ground state wavefunction. With this approximation, the NEB component of the $b$ singles cross section becomes
@@
For completeness we mention here an alternative formula proposed by Udagawa and Tamura Uda81 . Their final result can be again expressed in the form (74),
$\left.\frac{d^{2}\sigma}{dE_{b}d\Omega_{b}}\right|^{\mathrm{UT}}_{\mathrm{NEB}}=-\frac{2}{\hbar v_{i}}\rho_{b}(E_{b})\langle\varphi^{\mathrm{prior}}_{x}|W_{x}|\varphi^{\mathrm{prior}}_{x}\rangle,$ (81)
where $\varphi^{\mathrm{prior}}_{x}$ is the solution of a $x-A$ inhomogeneous equation similar to Eq. (79) but replacing in the source term the post-form transition operator, $V_{\mathrm{post}}$, by its prior form counterpart, $V_{\mathrm{prior}}\equiv U_{xA}+U_{bA}-U_{a}$, i.e.:
$(E^{+}_{x}-K_{x}-{U}_{x})\varphi^{\mathrm{prior}}_{x}({\vec{r}}_{x})=\langle{\vec{r}}_{x}\,\chi_{b}^{(-)}|V_{\mathrm{prior}}|\chi^{(+)}_{a}\phi_{a}\rangle.$ (82)
Although the final expressions for the IAV and UT models have the same formal structure, they lead to different predictions for the NEB cross sections. This is in contrast to the DWBA formula for transfer between bound states, where it is well known that the post and prior formulas are fully equivalent. This discrepancy led to a long-standing controversy between these two groups. At the heart of the discussion was the fact that the transformation of the post-form DWBA expression of IAV to its prior form gave rise to additional terms, not present in the UT prior formula. These additional terms guaranteed the post-prior equivalence for NEB, but they were regarded as unphysical by UT. Currently, there is consensus that that the IAV formula is the correct one, as corroborated by numerical comparisons to experimental data Jin15b .
@@
Figure 40: IAV calculations for the reaction 197Au(9Be,8Be*)198Au populating bound states of the residual nucleus. (a) Norm of the overlaps between the 9Be g.s. and the 8Be continuum pseudostates for different relative angular momenta $L$, as a function of the $\alpha$-$\alpha$ relative energy up to 20 MeV. Dashed lines are included as a guide. (b) Differential cross section, at $E_{\mathrm{cm}}=36$ MeV, as a function of the $\alpha$-$\alpha$ relative energy in the final 8Be system, for 0+ states (solid line) and $2^{+}$ states (dashed line). The symbols indicate the nominal energies of the g.s. ($0^{+}$) and the first $2^{+}_{1}$ resonance of 8Be. (c) Angle-integrated total transfer cross section as a function of $E_{\mathrm{cm}}$. The contribution from the $0^{+}$ (dot-dashed line) and $2^{+}$ (dotted) states of 8Be, together with their sum (blue solid), are compared to the experimental data of Ref. gollan21 . The IAV result rescaled by ab-initio Variational Monte Carlo (VMC) spectroscopic factors (dashed line) is also shown (see text). Adapted from Ref. Vil24 .
@@ 5.2 The Eikonal Hussein-McVoy formula (EHM)
    This particular choice has the virtue of taking into account breakup effects in the entrance channel. However, since each potential in Eq. (89) is evaluated in the corresponding fragment-target coordinate, the associated initial state wavefunction of the system would be a solution of a complicated three-body equation.
@@
With the particular choice of Eq. (89) and the assumption in Eq. (90) one obtains the following result for the $x$-channel wavefunction:
@@ 5.3 Interpretation of inclusive breakup data
In the previous sections, we have considered elastic breakup and nonelastic breakup (with the latter possibly including transfer and incomplete fusion) taking place in a breakup reaction. In actual inclusive breakup experiments these contributions will appear entangled in the data, although they produce some distinctive features. For example, the energy distribution of the observed fragment at a given scattering angle will exhibit a characteristic shape, as shown in Fig. 41 for a hypothetical $A(d,pX)$ reaction. To understand this spectrum, it is important to recall that a given proton energy and angle will univocally determine the excitation energy of the residual $n+A$ system. According to the proton energy (or, equivalently, the residual system excitation energy), we may distinguish the following regions:
Figure 41: Proton energy spectrum from a $A(d,pX)$ inclusive breakup reaction (blue line). The vertical dotted line marks the neutron separation threshold in the $A+n$ system. The low-energy peak arises from compound nucleus (CN) and pre-equilibrium (PE) processes followed by proton evaporation (magenta line).
@@
    As the excitation energy of the residual nucleus increases, so does the density of states and the low-lying bound states will merge to form a quasi-continuum. Here, the treatment with the aforementioned methods becomes more troublesome because of the impossibility of disentangling unambiguously the contribution of each state from the data. When one is interested in energy averaged cross sections, the combination of the IAV model of Sec. 5.1 with dispersive models provides an appealing alternative to the traditional methods.
@@
    As the excitation energy of the residual nucleus increases above the threshold, the narrow resonances disappear, the spectrum becomes continuous and structureless, giving rise to a bell-shaped bump. The most probable energy of the emerging proton can be obtained by assuming that the proton gets half the kinetic energy of the deuteron at the point of breakup, plus the Coulomb energy of the deuteron at that place, and half the internal energy of the deuteron ($E_{B}=-2.22$ MeV) (see Ref. Bla91 , p.509), i.e.,
@@
    In the former contributions, we have implicitly assumed that the observed particle (proton in this case) scatters elastically by the target nucleus. There will be situations in which this not the case; for example, when the projectile fuses completely with the target nucleus forming a compound nucleus that will eventually thermalize by emitting particles and gamma rays. Among these particles, there will be protons that will add up incoherently to the proton spectrum. These protons are typically emitted with low energy in the c.m. frame and will therefore contribute to the low energy part of the proton spectrum (see magenta line in Fig. 41). Some phenomena related to the fusion of weakly-bound nuclei will be discussed in the next section.
@@ 6 Fusion involving weakly bound nuclei
The previous sections have been devoted to the modeling of direct nuclear reactions. Fusion reactions involving weakly-bound nuclei display also distinctive features which require adequate reaction formalisms and considerations with respect to the case of well-bound nuclei. The topic of fusion with both stable and unstable nuclei has motivated many works and excellent review papers have been published in recent years, so we refer the reader to these works for a detailed account of the present status of the description of fusion (e.g. Can06 ; Can15 ; Can21 ). We shall discuss two phenomena which are the focus of many theoretical and experimental efforts by several groups. One is the phenomenon of complete fusion suppression. Complete fusion is conveniently defined as the process in which the whole charge of the projectile and target nuclei merge, giving rise to an excited compound nucleus that will subsequently decay by particle and/or gamma emission. The other phenomenon is the large observed yields compatible with the partial fusion of the projectile (incomplete fusion, ICF). Experimental results and theoretical calculations indicate that these two phenomena are actually related since they appear simultaneously and so a plausible explanation of the CF suppression might be in fact the leak of flux going to the ICF channels.
@@
Early analyses of these experiments tried to explain the phenomenon using coupled-channels calculations, including the coupling to low-lying excited states of the projectile and target Das99 ; Das02 ; Kum12 ; Zha14 ; Fan15 . Yet, these calculations systematically failed to reproduce the experimental suppression. This failure has been attributed to the omission in these calculations of the breakup of the projectile; a scenario was suggested in which the weakly bound projectile breaks up prior to reaching the fusion barrier, with the subsequent reduction of the complete fusion probability. This interpretation is supported by the presence of large $\alpha$ yields (in 6,7Li-induced reactions) as well as target-like residues which are consistent with the capture of one of the fragment constituents of the projectile, that is, ICF. To account for these observations, some authors have proposed a two-step scenario Das02 ; Dia07 consisting on the elastic dissociation of the projectile followed by the capture of one of the fragments by the target. However, calculations based on a three-dimensional classical dynamical model Dia07 , which incorporates this two-step breakup-fusion mechanism, can only explain a small fraction of the observed CF suppression for 9Be Coo16 and 8Li Coo18 reactions. More encouraging results have been obtained with different methods based on the CDCC formalism, as described in the following subsections.